### v5

@@ 2.2 Illustrative Example
$\footnotesize P(F)=$ $\sum\limits_{A,B,C,D,E,G}P(A)P(B)P(C)P(D|A,B)P(E|C)P(F|A)P(G|D,E)$ (4)
$=$ $\sum\limits_{A}P(F|A)P(A)\left(\sum\limits_{D,E,G}P(G|D,E)\left(\sum\limits_{B}P(D|A,B)P(B)\left(\sum\limits_{C}P(E|C)P(C)\right)\right)\right)$ (5)
@@
Contrast this with the cluster mapping shown through shading (e.g. $A$ and $F$ are in the same cluster, as are $D$ and $G$; all others in a cluster of their own) in Fig. 1. We illustrate clustering with a simple approach where probabilities from an antecedent cluster are multiplied with those in the current cluster then marginalised using an arbitrary ordering. The resultant computation (5) has maximum dimensionality of $4$ due to interspersed multiplication and marginalisation; recall marginalisation happened at the end in (4). In this example, clustering greatly reduces computation and memory cost. It also demonstrates cost dependency, since the computation cost required at a descendant cluster is dependent on the mapping and outcomes of antecedent clusters. Finally, it demonstrates the forwards and backwards nature of inference, such as forwards propagation from cluster $E$ to cluster $\{D,G\}$, and backwards propagation from $\{D,G\}$ to cluster $\{A,F\}$.
@@
Figure 1: Example DAG showing nodes, directed edges, and layers where shading denotes clusters obtained using DCMAP. Assume all nodes have binary states.
2.3 Overview and Notation
@@
[…]
@@
$c()$ Transition cost for one step of dynamic programming; a function of a cluster-layer and its child cluster-layer(s) (Section 2.3.3).
@@
$kl$, $\boldsymbol{X}_{kl}$ Cluster-layer label and set of cluster-layer nodes, respectively (Section 2.3.3).
@@ 2.3.1 Layers
We assume that $H$ is fully connected, since cluster mapping could be applied independently to each connected component. We use a simple breadth-first search [16] to compute $l(\boldsymbol{X})$ in $O(n)$ time (Algorithm 1).
@@
2.3.2 Clusters
Definition 2.2.
Nodes are mapped to clusters using integer labels $u(X)=k$ where $1\leq u(X)\leq m$ for $m$ unique clusters, $m\leq n$. Denote the set of nodes in cluster $k$ as $\boldsymbol{X}_{k}$. Denote $u^{\prime}(X)$ as a proposed cluster mapping for node $X$. Clusters must be contiguous (Definition 2.3).
Definition 2.3.
A cluster $\boldsymbol{X}_{k}$ must be contiguous such that, given any two nodes $X_{1},X_{2}\in\boldsymbol{X}_{k}$, there are no parent-child paths $\boldsymbol{X}_{p}=\{X_{1},\ldots,X^{\prime},\ldots,X_{2}\}$ with a node $X^{\prime}$ mapped to a different cluster.
As an example, if in Fig. 1 nodes $B,G$ were in cluster $1$ and $D$ in cluster $2$, then that breaks the contiguity of cluster $1$ and introduces a cycle between cluster $1$ and $2$ via the path $\{B,D,G\}$. If instead nodes $D,G$ were in cluster $1$ and $B$ in cluster $2$, then contiguity is maintained and there are no cycles between cluster $1$ and $2$.
Definition 2.4.
@@ 2.3.2 Clusters
This definition holds because by Definition 2.3, a contiguous cluster cannot have any gap layers, i.e. $l^{0}_{k},\ldots,l^{1}_{k}$ is a counting sequence. Any gap layer $l^{\prime}:l^{0}_{k}\leq l^{\prime}\leq l^{1}_{k}$ implies the existence of at least one path from $\boldsymbol{X}_{kl^{0}_{k}}$ through a different cluster-layer $\boldsymbol{X}_{k^{\prime}l^{\prime}}$ to $\boldsymbol{X}_{kl^{1}_{k}}$ which violates contiguity, by definition of a layer as the length of the longest path (Definition 2.1).
Definition 2.5.
@@
2.3.3 Dynamic Programming
We use dynamic programming to find optimal cluster mappings. Generally, dynamic programming is formulated as the optimisation of an objective function (or cost function) $G$ over a series of decisions or actions $u_{k}$ for $k=0,\ldots,K-1$ where the optimal cost $G^{\ast}$ over an interval $[k,K]$ is [3]:
$G_{(k,K)}^{\ast}(x)=\inf_{u_{k}\in\mathcal{A}_{k}}\left[G_{(k,K)}\left(x_{k},u_{k},U_{(k+1,K-1)}^{\ast}\right)\right]$ (6)
where $U_{(k+1,K-1)}^{\ast}$ is the optimal sequence of actions $u_{k+1},\ldots,u_{K-1}$ for the interval $[k+1,K-1]$, $\mathcal{A}_{k}$ is the set of possible actions at stage $k$, $x$ is the state at stage $K$ as a consequence of optimal choice of $u_{k}$ at state $x_{k}$, and $\inf$ is the infimum operator which returns the argument that gives the minimum expression. Intuitively, given the current state $x_{k}$ and an optimal sequence of actions at stages $k+1$ to $K-1$ leading to state $x=x_{K}$, we choose the action $u_{k}$ to take at the current stage to minimise the overall cost in going from stage $k$ to $K$. This formulation is the most flexible as the cost function can vary (i.e time-varying in the literature, but we refer to as stage varying to avoid confusion with time in a DBN) across intervals $[k,K]$, and does not have to be additive. This is useful as there can be multiple clusters in a given layer of the DAG to cluster (Section 2.4). However, computational complexity grows in NP time with graph size and with the number of proposals $|\mathcal{A}_{k}|$ at each stage [16].
@@ 2.3.3 Dynamic Programming
$G^{\ast}_{k+1}(x)=\inf_{u\in\mathcal{A}}\left[c(x,u)+G_{k}^{\ast}\left(f(\boldsymbol{x},\boldsymbol{u})\right)\right]$ (7)
where $f(\boldsymbol{x},\boldsymbol{u})$ is a sequence of states and actions, starting at stage $k=0$, resulting in state $x$ at stage $k$ with optimal cost $G^{\ast}_{k}$, and $c$ is the transition cost of enacting action $u$ at state $x$. An additive and time-invariant system greatly reduces computational complexity. Optimisation algorithms are built around the objective function, and Equation (7) is the basis for many heuristic search techniques such as A*. A heuristic estimate of the remaining cost over stages $[k+1,K]$ is used in A* to prioritise the search and potentially reduce computation time, such as for BN learning [18].
2.4 High-Level Algorithm
@@
Initialisations (Section 3.1.1) Pop $kl$ from $Q^{\prime}$, or activate a branch if no eligible $kl$ and try again (Section 3.1.2) Prune dominated branches (linked or global threshold $G_{\text{min}}$) (Section 3.1.3) Propose cluster mappings for $\boldsymbol{X}_{kl}$ based on $\hat{\textrm{U}}$ (Section 3.1.4) Compute costs and update branches (Section 3.1.5). Least cost branch is active. If a layer is completely mapped (Section 3.1.6), update linked branches, update $G_{\text{min}}$ if $l=l_{\text{max}}$. Update possible $kl$ mappings for this branch in $\hat{\textrm{U}}$ and put this $kl$ and branch into $Q$ (Section 3.1.7) For each $\boldsymbol{X}_{kl}$ proposalWhile $Q^{\prime}$ is not empty
Figure 2: Example network showing nodes, directed edges, and layers where shading denotes clusters obtained using DCMAP. Assume all nodes have binary states.
2.5 Objective Function and Heuristic
@@
$L=\max\Bigl{(}l\bigl{(}Par(\boldsymbol{X}_{l})\bigr{)}\Bigr{)}$ (8)
Substituting into the general objective function of Bellman and Dreyfus [3] (6), the objective function over layers $[l,L]$ is:
$G_{(l,L)}^{\ast}(x)$ $=\inf_{u(\boldsymbol{X}_{l})\in U_{l}}\left[G_{(l,L)}\left(\boldsymbol{X}_{l},u(\boldsymbol{X}_{l}),U_{(l+1,L)}^{\ast}\right)\right]$ (9)
@@
As disjoint nodes are conditionally independent and nodes in a layer or cluster-layer are disjoint by Definition 2.4, we assume that the objective function cost of disjoint cluster-layers are independent and thus additive. Moreover, by definition of $L$ (8) as the latest parent layers given nodes at layer $l$, an additive and time-invariant formulation for the objective function could be obtained by finding the lowest cost cluster-mapping (the infimum) over all intermediary cluster-layers. This replaces the $U^{*}_{l+1,L}$ term in (6) and makes $G$ time invariant. Thus, (9) can be re-written as:
@@
$\inf_{u(\boldsymbol{X}_{l+1,\ldots,L})}\sum_{l^{\prime}=l+1}^{L}{\sum_{k=1}^{m_{l^{\prime}}}{c\Bigl{(}\boldsymbol{X}_{kl^{\prime}},u(\boldsymbol{X}_{kl^{\prime}}),Child(\boldsymbol{X}_{kl^{\prime}}),u\bigl{(}Child(\boldsymbol{X}_{kl^{\prime}})\bigr{)}\Bigr{)}}}$ (10)
where the optimal cost $G^{*}(\boldsymbol{X}_{L})$ is only a function of the optimal cost at layer $l$ and cluster-layer mappings between $l+1$ and $L$ inclusive, notated as $u(\boldsymbol{X}_{l+1,\ldots,L})$, and $c$, a positive, non-zero cost associated with a cluster-layer $\boldsymbol{X}_{kl^{\prime}}$ and its cluster mapping $u(\boldsymbol{X}_{kl^{\prime}})$. Note that $m_{l^{\prime}}$ is the number of unique clusters at layer $l^{\prime}$. In the BN inference context (Section 2.2 and (3), cost dependency arises from propagation between child cluster-layers $Child(\boldsymbol{X}_{kl^{\prime}})$, their cluster mapping, and a parent cluster-layer $kl^{\prime}$ and its cluster mapping; potentially, they could be of the same cluster.
Equation (10) provides the basis for our algorithm. However, in algorithms like A* [16], potential actions (i.e. cluster mappings) are evaluated prospectively using the objective function at each iteration. This rapidly becomes infeasible in our cluster mapping context due to the combinatorial explosion of mappings over multiple layers $u(\boldsymbol{X}_{l+1,\ldots,L})$. One approach to overcome this challenge is to instead apply our objective function (10) retrospectively after potentially multiple iterations of the algorithm, to prune dominated branches (Section 3.2 and Lemma 3.2).
Definition 2.6.
@@
$h\Bigl{(}\boldsymbol{X}\setminus\bigl{(}\boldsymbol{X}_{kl:k\in M_{l}}\cup\boldsymbol{X}_{0..l-1}\bigr{)}\Bigr{)}$ (11)
where $M_{l}$ is a set of clusters at layer $l$ that have been updated, $\setminus$ is set difference, and $h$ is a heuristic cost that is an upper bound for all the remaining, non-updated nodes. We refer to nodes for which the cost has been calculated as updated or popped. Potentially, there may be some nodes remaining in layer $l$ that have not been updated which are estimated as part of the $h$ term (11). Note that it is only possible to update cluster-layers at layer $l$ when all nodes in layer $l-1$ and earlier layers have been updated at least once.
2.6 Search Space Size
Using the concept of layers, we can approximate the size of the search space of our cluster mapping problem. Naively, a DAG with $n$ nodes could be mapped to $k=1..n$ clusters thus resulting in $n^{n}$ possible mappings. However, due to the requirement for cluster contiguity (Definition 2.3), there are many infeasible proposals. A simple approach to ensuring contiguity is to either map a node to a new cluster unique to itself, or to an existing cluster of a child node (Section 3.1.4). The overall search space size $\eta$ arises from the combination of layer by layer cluster mapping combinations, i.e., the product of products (12).
$\eta=\prod_{l=1..l_{\text{max}}}{\prod_{X_{i}\in\boldsymbol{X}_{l}}{\bigl{(}\left|Child(X_{i})\right|+1\bigr{)}}}$ (12)
As might be expected, the search space grows exponentially with layers, and the out-degree of the DAG can also significantly grow the size of the search space [4]. In- and out-degree refer to the number of parent or child nodes, respectively [4]. For instance, the case study DBN [26], with small world connectivity, has just 25 nodes but a search space of 9.9 billion mapping combinations. This study and corresponding DCMAP solution are described in Section 4.
3 Algorithm and Analysis

### v6

@@ 2.2. Illustrative Example
(4) $\footnotesize P(F)=$ $\sum\limits_{A,B,C,D,E,G}P(A)P(B)P(C)P(D|A,B)P(E|C)P(F|A)P(G|D,E)$
(5) $=$ $\sum\limits_{A}P(F|A)P(A)\left(\sum\limits_{D,E,G}P(G|D,E)\left(\sum\limits_{B}P(D|A,B)P(B)\left(\sum\limits_{C}P(E|C)P(C)\right)\right)\right)$
@@
Contrast this with the cluster mapping shown through shading (e.g. $A$ and $F$ are in the same cluster, as are $D$ and $G$; all others in a cluster of their own) in Fig. 1. We illustrate clustering with a simple approach where probabilities from an antecedent cluster are multiplied with those in the current cluster then marginalised using an arbitrary ordering. The resultant computation (5) has maximum dimensionality of $4$ due to interspersed multiplication and marginalisation; recall marginalisation happened at the end in (4). In this example, clustering greatly reduces computation and memory cost. It also demonstrates cost dependency, since the computation cost required at a descendant cluster is dependent on the mapping and outcomes of antecedent clusters. Finally, it demonstrates the forwards and backwards nature of inference, such as forwards propagation from cluster $E$ to cluster $\{D,G\}$, and backwards propagation from $\{D,G\}$ to cluster $\{A,F\}$. For the full example, see Appendix A.
@@
Figure 1. Example DAG showing nodes, directed edges, and layers where shading denotes clusters obtained using DCMAP. Assume all nodes have binary states.
2.3. Overview and Notation
@@
Table 1. Summary of notation. Generally, we use lower case notation (e.g. $g,u$) for local calculation, upper case (e.g. $G,U$) for global search and optimisation, non-italicised (e.g. $\text{G},\text{U}$) for data structures in our algorithm, and bold for a set. We notate $k=1..n$ to denote k=1,2,,nk=1,2,...,n for brevity.
@@
$c()$ Transition cost for one step of dynamic programming (10); positive, non-zero function of nodes and associated conditional probabilities in a cluster-layer and child cluster-layer(s).
@@
klkl, $\boldsymbol{X}_{kl}$ Cluster-layer label and set of cluster-layer nodes, respectively (Section 2.3.3).
@@ 2.3.1. Layers
We assume that $H$ is fully connected, since cluster mapping could be applied independently to each connected component. We use a simple breadth-first search (Lavalle2006) to compute $l(\boldsymbol{X})$ in $O(n)$ time (Algorithm 1).
@@
2.3.2. Clusters
Definition 2.
Nodes are mapped to clusters using integer labels $u(X)=k$ where $1\leq u(X)\leq m$ for $m$ unique clusters, $m\leq n$. Denote the set of nodes in cluster $k$ as $\boldsymbol{X}_{k}$. Denote $u^{\prime}(X)$ as a proposed cluster mapping for node $X$.
Clusters must be contiguous (Definition 3) to avoid possible introduction of cycles between clusters.
Definition 3.
A cluster $\boldsymbol{X}_{k}$ is contiguous iff given any two nodes $X_{1},X_{2}\in\boldsymbol{X}_{k}$, there are no parent-child paths $\boldsymbol{X}_{p}=\{X_{1},\ldots,X^{\prime},\ldots,X_{2}\}$ with a node $X^{\prime}$ mapped to a different cluster.
BN inference assumes a DAG with no cycles to ensure convergence when propagating evidence (Pearl1988). This extends to clusters (or cliques) to enable computationally efficient propagation of evidence between clusters as part of BN inference (Lauritzen1988). As an example, if in Fig. 1 nodes $B,G$ were in cluster $1$ and $D$ in cluster $2$, then that breaks the contiguity of cluster $1$ and introduces a cycle between cluster $1$ and $2$ via the path $\{B,D,G\}$. If instead nodes $D,G$ were in cluster $1$ and $B$ in cluster $2$, then contiguity is maintained and there are no cycles between cluster $1$ and $2$. Since the underlying DAG is acyclic, it is easy to see that, with contiguity, the resulting clusters are also acyclic.
Definition 4.
@@ 2.3.2. Clusters
Definition 5.
@@
2.3.3. Dynamic Programming
We use dynamic programming to find optimal cluster mappings. Generally, dynamic programming is formulated as the optimisation of an objective function (or cost function) $G$ over a series of decisions or actions $u_{k}$ for $k=0,\ldots,K-1$ where the optimal cost $G^{\ast}$ over an interval $[k,K]$ is (Bellman1962):
(6) $G_{(k,K)}^{\ast}(x)=\inf_{u_{k}\in\mathcal{A}_{k}}\left[G_{(k,K)}\left(x_{k},u_{k},U_{(k+1,K-1)}^{\ast}\right)\right]$
where $U_{(k+1,K-1)}^{\ast}$ is the optimal sequence of actions $u_{k+1},\ldots,u_{K-1}$ for the interval [k+1,K1][k+1,K-1], $\mathcal{A}_{k}$ is the set of possible actions at stage $k$, $x$ is the state at stage $K$ as a consequence of optimal choice of $u_{k}$ at state $x_{k}$, and $\inf$ is the infimum operator which returns the argument that gives the minimum expression. Intuitively, given the current state $x_{k}$ and an optimal sequence of actions at stages $k+1$ to K1K-1 leading to state $x=x_{K}$, we choose the action $u_{k}$ to take at the current stage to minimise the overall cost in going from stage $k$ to $K$. This formulation is the most flexible as the cost function can vary (i.e time-varying in the literature, but we refer to as stage varying to avoid confusion with time in a DBN) across intervals $[k,K]$, and does not have to be additive. This is useful as there can be multiple clusters in a given layer of the DAG to cluster (Section 2.4). However, computational complexity grows in NP time with graph size and with the number of proposals $|\mathcal{A}_{k}|$ at each stage (Lavalle2006).
@@ 2.3.3. Dynamic Programming
(7) $G^{\ast}_{k+1}(x)=\inf_{u\in\mathcal{A}}\left[c(x,u)+G_{k}^{\ast}\left(f(\boldsymbol{x},\boldsymbol{u})\right)\right]$
where $f(\boldsymbol{x},\boldsymbol{u})$ is a sequence of states and actions, starting at stage $k=0$, resulting in state $x$ at stage $k$ with optimal cost $G^{\ast}_{k}$, and $c$ is the transition cost of enacting action $u$ at state $x$. An additive and time-invariant system greatly reduces computational complexity. Optimisation algorithms are built around the objective function, and Equation (7) is the basis for many heuristic search techniques such as A*. A heuristic estimate of the remaining cost over stages $[k+1,K]$ is used in A* to prioritise the search and potentially reduce computation time, such as for BN learning (Lu2021).
2.4. High-Level Algorithm
@@
Initialisations (Section 3.1.1) Pop klkl from $Q^{\prime}$, or activate a branch if no eligible klkl and try again (Section 3.1.2) Prune dominated branches (linked or global threshold $G_{\text{min}}$) (Section 3.1.3) Propose cluster mappings for $\boldsymbol{X}_{kl}$ based on $\hat{\textrm{U}}$ (Section 3.1.4) Compute costs and update branches (Section 3.1.5). Least cost branch is active. If a layer is completely mapped (Section 3.1.6), update linked branches, update $G_{\text{min}}$ if $l=l_{\text{max}}$. Update possible klkl mappings for this branch in $\hat{\textrm{U}}$ and put this $kl$ and branch into $Q$ (Section 3.1.7) For each $\boldsymbol{X}_{kl}$ proposal While $Q^{\prime}$ is not empty
Figure 2. Illustration of the high level components of the proposed algorithm DCMAP.
2.5. Objective Function and Heuristic
@@
(8) $L=\max\Bigl(l\bigl(Par(\boldsymbol{X}_{l})\bigr)\Bigr)$
Substituting into the general objective function of (Bellman1962) (6), the objective function over layers $[l,L]$ is:
(9) $G_{(l,L)}^{\ast}(x)$ $=\inf_{u(\boldsymbol{X}_{l})\in U_{l}}\left[G_{(l,L)}\left(\boldsymbol{X}_{l},u(\boldsymbol{X}_{l}),U_{(l+1,L)}^{\ast}\right)\right]$
@@
As disjoint nodes are conditionally independent and nodes in a layer or cluster-layer are disjoint by Definition 4, we assume that the objective function cost of disjoint cluster-layers are independent and thus additive. Moreover, by definition of $L$ (8) as the latest parent layers given nodes at layer $l$, an additive and time-invariant formulation for the objective function could be obtained by finding the lowest cost cluster-mapping (the infimum) over all intermediary cluster-layers. This replaces the $U^{*}_{l+1,L}$ term in (6) and makes $G$ time invariant. Thus, (9) can be re-written as:
@@
(10) $\inf_{u(\boldsymbol{X}_{l+1,\ldots,L})}\sum_{l^{\prime}=l+1}^{L}{\sum_{k=1}^{m_{l^{\prime}}}{c\Bigl(\boldsymbol{X}_{kl^{\prime}},u(\boldsymbol{X}_{kl^{\prime}}),Child(\boldsymbol{X}_{kl^{\prime}}),u\bigl(Child(\boldsymbol{X}_{kl^{\prime}})\bigr)\Bigr)}}$
where the optimal cost $G^{*}(\boldsymbol{X}_{L})$ is only a function of the optimal cost at layer $l$ and cluster-layer mappings between $l+1$ and $L$ inclusive, notated as $u(\boldsymbol{X}_{l+1,\ldots,L})$, and $c$ is a positive, non-zero cost that is a function of nodes, cluster mappings and implicitly associated conditional probabilities in a cluster-layer and child cluster-layer(s). Note that $m_{l^{\prime}}$ is the number of unique clusters at layer $l^{\prime}$. In the BN inference context (Section 2.2 and (3), cost dependency arises from propagation between child cluster-layers $Child(\boldsymbol{X}_{kl^{\prime}})$, their cluster mapping, and a parent cluster-layer $kl^{\prime}$ and its cluster mapping; potentially, they could be of the same cluster.
Equation (10) provides the basis for our algorithm. However, in algorithms like A* (Lavalle2006), potential actions (i.e. cluster mappings) are evaluated prospectively using the objective function at each iteration. This rapidly becomes infeasible in our cluster mapping context due to the combinatorial explosion of mappings over multiple layers $u(\boldsymbol{X}_{l+1,\ldots,L})$. One approach to overcome this challenge is to instead apply our objective function (10) retrospectively after potentially multiple iterations of the algorithm, to prune dominated branches (Section 3.2 and Lemma 3.2).
Definition 6.
@@
(11) $h\Bigl(\boldsymbol{X}\setminus\bigl(\boldsymbol{X}_{kl:k\in M_{l}}\cup\boldsymbol{X}_{0..l-1}\bigr)\Bigr)$
where $M_{l}$ is a set of clusters at layer $l$ that have been updated, $\setminus$ is set difference, and $h$ is a heuristic cost that is an upper bound for all the remaining, non-updated nodes. We refer to nodes for which the cost has been calculated as updated or popped. Potentially, there may be some nodes remaining in layer $l$ that have not been updated which are estimated as part of the $h$ term (11). Note that it is only possible to update cluster-layers at layer $l$ when all nodes in layer l1l-1 and earlier layers have been updated at least once.
2.6. Search Space Size
Using the concept of layers, we can approximate the size of the search space of our cluster mapping problem. Naively, a DAG with $n$ nodes could be mapped to $k=1..n$ clusters thus resulting in $n^{n}$ possible mappings. However, due to the requirement for cluster contiguity (Definition 3), there are many infeasible proposals. A simple approach to ensuring contiguity is to either map a node to a new cluster unique to itself, or to an existing cluster of a child node (Section 3.1.4). The overall search space size $\eta$ arises from the combination of layer by layer cluster mapping combinations, i.e., the product of products (12).
(12) $\eta=\prod_{l=1..l_{\text{max}}}{\prod_{X_{i}\in\boldsymbol{X}_{l}}{\bigl(\left|Child(X_{i})\right|+1\bigr)}}$
As might be expected, the search space grows exponentially with layers, and the out-degree of the DAG can also significantly grow the size of the search space (Buluc2016). In- and out-degree refer to the number of parent or child nodes, respectively (Buluc2016). For instance, the case study DBN (Wu2018), with small world connectivity, has just 25 nodes but a search space of 9.9 billion mapping combinations. This study and corresponding DCMAP solution are described in Section 4.
3. Algorithm and Analysis